# Attention as a Hypernetwork

### v3

@@ 3.1 Compositional generalization
We first investigate how well multi-head attention based transformer models are able to learn our fuzzy logic tasks in-context and compositionally generalize to held out tasks. We vary the number of examples presented in context as well as the fraction of combinations of fuzzy logic terms held out from the training set. Figure 2C compares standard multi-head softmax attention and linear attention as well as our HYLA model in terms of their coefficient of determination ($R^{2}$) on the queries of held out tasks. We find that all models are able to solve the fuzzy logic task given sufficient examples are presented in context. As the fraction of held out tasks increases, compositional generalization performance starts decreasing for all models. Interestingly, HYLA is less affected by this decline than softmax and linear attention. We hypothesize that the more expressive value network allows to more efficiently represent the nonlinear terms being composed and thereby strengthens the inductive bias towards compositionality. Consistent with this interpretation, in-distribution the models show comparably small differences in performance (see A2). In Appendix A.1 we conduct additional model ablations that separate the contributions of the RMSHead normalization and the nonlinearity added to the value network, showing that both components in combination are required to obtain the observed performance improvements.
Our compositional split only contains novel combinations of known terms. A priori, it might be possible that the system learns to compose the basic fuzzy logic operations in a way that allows it to generalize to any fuzzy logic function, including those that contain novel combinations of unknown terms. However, testing on functions obtained as novel combinations of $K$ unknown terms, we find that none of the models considered here is able to solve such a task (see Figure A4 in the appendix).
@@ 3.2 Latent code structure
Next, we analyze the structure of the latent code for the different models solving the task. As suggested by Eq. 4, the attention scores for a given key-query pair across the head dimension configure the computation performed in the value network and accordingly we would expect it to reflect the fuzzy logic function that needs to be implemented to correctly predict the output of the query token. To test this hypothesis, we collect the attention scores for the final query token attending to itself, obtaining a vector of dimension $H$ for each layer and task. For this purpose we use the tasks held out from training, varying the query token while keeping inputs shown in-context fixed. In order to visualize the $H$-dimensional points we thereby obtain for each layer, we reduce the dimensionality using tSNE [27]. The results of this analysis are shown in Figure 2D (for an extended figure showing all layers see Figure A3). In line with our hypothesis, the latent codes for each model form clusters. The structure of these clusters is only partially explained by the target value of the function for a given query value. Strikingly, coloring data points according to the fuzzy logic terms present underlying each function reveals that clusters in the latent codes closely correspond to the functions. Indeed, it is possible to decode the terms underlying each function from the latent code using a logistic regression classifier which is trained to predict the function terms given the latents and is evaluated on held out tasks as shown in Figure 2B.
@@
Figure 3: sraven. A Illustration of sraven task generation and the construction of the compositional generalization split. B Example problem instance illustrating a key challenge of the original Raven Progressive Matrices of finding correspondences (adapted from [28]). When attempting to solve this instance, different hypotheses over which figural elements are governed by a consistent rule across rows are possible. This is akin to different orderings of the symbolic features.
Next we will introduce an abstract reasoning task based on Raven Progressive Matrices [3] we refer to as sraven. We develop this task to provide a challenging benchmark that specifically probes symbolic reasoning capabilities while giving us fine-grained control over the constituent parts of each task in order to assess compositional abilities and analyze latent code structure.
4.1 Abstract reasoning based on symbolic Raven Progressive Matrices
The original Raven Progressive Matrices test is a classic human intelligence test that probes the ability to induce abstract relations [3]. It requires test subjects to generate and verify a potentially large number of hypotheses over rules that can parsimoniously explain a given problem instance [28]. While there have been previous machine learning datasets inspired by Raven Progressive Matrices [29, 30, 31, 32], we seek to study a setting that is both symbolic, compositional and models an aspect which to the best of our knowledge has not been explicitly considered in prior benchmarks: we seek to model the key difficulty of Raven tasks referred to as finding correspondences by Carpenter et al. [28] which requires to search through a large number of possible hypotheses. We discuss existing Raven-based benchmarks in Section 6.
@@ 4.2 Task description
Figure 3A illustrates the generative process of sraven. For each task, a matrix of eight context panels is presented and the goal is to predict the contents of the final query panel. In the original Raven Progressive Matrices Test, each panel typically displays a combination of geometrical shapes as for instance shown in Figure 3B. For our symbolic version, each panel is defined as a vector of integers of length $M$ (in our experiments we set $M=4$ unless stated otherwise). For each dimension $m\in\{1,2,\ldots,M\}$ one of $R=8$ possible rules is sampled as well as three corresponding sequences of three integers generated by the respective rule (see Figure 3A and Appendix B.3 for more details on the rules). We restrict rules to apply within each row while generally rules could also be applied both across columns and across rows. As a result we obtain nine panels where each row can be explained by the same set of underlying rules. In order to maintain a finite set of possible symbols both for the context panels and the query panel, we limit all tasks to contain only integers $\{0,1,2,\ldots,K-1\}$ (in our experiments we set $K=8$). Arithmetic rules such as progression, sum or difference are therefore defined using modular arithmetic modulo $K$.
We use the first eight panels as context panels and present them sequentially to a sequence model which needs to predict each dimension of the final query panel independently. A prediction is considered correct if and only if all subpredictions were correct. In the original Raven task, a number of possible answer panels are presented alongside the context panels and test subjects are asked to pick the one which contains the correct answer [3]. Since in our case the correct answer will consist of a combination of a finite set of known symbols, we can ask the agent solving the task to directly predict the missing symbols. Conveniently this helps avoid potential shortcuts for solving a task that might arise from a biased generative process to create multiple-choice answers as a popular prior Raven-based benchmark has been found to be affected by [30, 32].
One of the key features of sraven is its compositional structure. Each task consists of a combination of a finite set of rule combinations. To systematically test whether a system trained on a subset of rule combinations can generalize to unseen combinations, we split all rule combinations into a train set and a test set. Unless noted otherwise, we hold out $25\%$ of all possible rule combinations for evaluation in our experiments. Note that permutations of rule combinations (e.g. AB and BA) are considered to be the same and will only appear in one of the two sets. This ensures that at test time the agent is confronted with instances of tasks whose underlying combination of rules it has never encountered during training.
@@ 4.3 Finding correspondences
Making the task symbolic instead of presenting images of geometric shapes comes with the advantage of a smaller input space allowing us to scale to larger model and data sizes in our experiments. However, it removes the visual representation learning step of discovering the right symbolic abstractions. We argue that while this is a difficult problem worth studying by itself, it is typically not the main source of difficulty for the original Raven tasks which mostly consist of simple visual features such as numerosity, shape, or position. As Carpenter et al. [28] argue in their theoretical account of the original Raven test, what makes many tasks difficult is the problem of finding correspondences of rules to feature dimensions.
This is best illustrated when considering the example in Figure 3B showing an adapted version of Figure 5 of [28]. When first attempting to solve this task, a test subject might hypothesize that there is a rule guiding the numerosity and orientations per shape. For example one might to try find a rule that governs all curved shapes, one rule that governs all lines and one rule that governs all rectangles. Translated into a symbolic encoding, this is akin to ordering symbols by the shape and trying to find rules along each dimension in this ordering. However, this approach turns out to not provide a parsimonious answer. The correct answer can instead be obtained by grouping the objects by their orientations, i.e. by finding a rule that applies to all horizontal objects and a rule that applies to all vertical objects. In the symbolic encoding this new correspondence of features amounts to a new ordering of the same symbolic encoding where symbols are now ordered by orientation and one needs to find rules guiding the numerosity and shape per orientation.
To model the property of finding correspondences in sraven we similarly allow for permutations of the features. Specifically, we sample and apply a column-specific permutation of the features when generating each task such that the entries of the vectors encoding each panel are permuted with a consistent permutation for a given column across all rows. This makes finding the correct solution difficult since the number of possible hypotheses grows significantly (see Appendix B.1 for more details). Note that in principle a problem instance might have multiple valid but divergent answers making the task ill-posed. In the settings considered here this happens for less than half a percent of the problem instances. See Appendix B.2 for more details on handling such ambiguities.
@@ 4.4 Results
Figure 5: Latent code structure of sraven. A tSNE visualizations of the final layer latent codes for the final query token colored by the magnitude of the predicted target value. B Same as A but colored by the ground-truth rule the model needs to apply to the query token to generated the correct prediction. C OOD accuracy for varying numbers of attention heads. For a single head the hypernetwork mechanism is absent which hampers OOD generalization. D The difficulty of sraven can be parametrically controlled by varying the number of modules $M$. E Heatmap showing the pairwise cosine similarity between the average latent code of each rule for HYLA revealing how semantically related rules form clusters. For instance rule F (sum) and G (difference) are implemented with a very similar latent code indicating that the same code might be reused by flipping the sign of the operands. F Decoding performance of a logistic regression classifier trained to predict the ground-truth sraven rule based on the latent code at the final query token of training tasks and evaluated on unseen OOD tasks revealing that the latent code is predictive of the implemented rule.
@@
We first evaluate to what extent transformers can compositionally generalize on sraven as we increase the number of problem instances trained on and scale up the models both in terms of width and depth. Figure 4 shows the results of these scaling experiments. Given sufficient data and model size, all model classes are able to solve 80% of problem instances created from tasks held out during training. The inductive bias of making the value network more expressive by introducing a nonlinearity in HYLA seems to be beneficial for this task. Especially when presented with few problem instances and for small model sizes it outperforms both linear and softmax attention. In contrast, disrupting the hypernetwork mechanism by decreasing the number of heads to $H=1$ noticeably decreases performance as shown in Figure 5C.
@@
We next analyze the latent code for the different models solving the task. To this end we collect the attention scores of the final query token attending to itself. Following the perspective of attention as a hypernetwork, we expect latent codes corresponding to the same rule to form clusters. Indeed, Figure 5B reveals a strikingly structured code for the final layer across models with clusters closely matching the rule to be inferred to correctly predict the query output for each given task. In comparison, Figure 5A colors the points in latent space by the magnitude of the inputs the query needs to attend to make the correct prediction. This tends to explain clusters more prominently in early layers as shown in the appendix in Figure A5 and for a 16 layer softmax transformer in Figure A6.
To obtain a better understanding of the semantics of the latent code, we explore the pairwise cosine similarity between average latent codes for each rule across tasks for HYLA. Figure 5C shows that semantically related rules like sum/difference can form clusters, suggesting that latent codes might even be reused to solve related task operations. Finally, we train a logistic regression classifier to predict the ground-truth sraven rule based on the latent code at the final query token of in-distribution sraven tasks and evaluate it on unseen OOD tasks. This reveals that especially for later layers the latent code is strongly predictive of the subfunctions performed by the network.
Figure 6: Language modeling. Performance of autoregressively trained decoder-only transformer models with 50M parameters over the course of training on the C4 dataset with 130B tokens [33].
@@ 5 Language modeling
Multi-head attention has proven especially effective for language modeling. With language itself being highly compositional, it stands to question whether the hypernetwork mechanism inside of multi-head attention might play a part in explaining its success. To shed light on this question, we conducted language modeling experiments and evaluate the performance of HYLA in comparison to linear and softmax attention. We train decoder-only transformer models with 50M parameters autoregressively for 130 Billion tokens on the C4 dataset [33]. We find that strengthening the hypernetwork mechanism via HYLA improves performance over linear attention and performs closely to softmax attention despite being a linear attention variant itself. This is noteworthy considering the importance of softmax in language modeling hypothesized to be due to its role in binding and associative recall problems on which linear attention methods typically struggle [34, 35, 13]. The fact that reinforcing the hypernetwork mechanism inside of multi-head attention as done by HYLA helps close the gap of linear attention to softmax attention suggests that the hypernetwork mechanism might be of practical relevance for understanding large-scale transformer models.
@@ 6 Related work
Prior work that has studied the question why the attention mechanism benefits from multiple heads has counterintuitively found that it is possible to prune almost all but one head in some layers after training, sacrificing only relatively small amounts of performance [36, 37]. It has therefore been speculated that equipping the attention mechanism with multiple heads primarily aids stability during training [38]. The hypernetwork perspective of attention offers an alternative account, suggesting that multiple heads create a way of configuring compositional computations in the value network. Seen from this perspective, the observation of singular heads dominating could point to a connection to the phenomenon of module collapse often observed when training modular systems in practice [39, 40, 41, 42].
@@
Consistent with our observation on scaling model size on the sraven task, [18] find that as pretrained large language models are scaled, their ability to compositionally generalize on in-context semantic parsing tasks improves. Similarly, outside the in-context learning regime, [43] have demonstrated that pretrained large language models outperform many architectures specialized towards compositional generalization. Still, to what extent the ability for compositional generalization extends beyond the training distribution remains openly debated [19, 20, 21].
@@
A number of Raven-inspired tasks have been introduced to assess abstract reasoning capabilities of neural networks [29, 30, 31, 32]. Common to all these variants including our version is an underlying generative task model based on a finite number of possible rules that are applied on one or more of the feature dimensions for each panel. Different from our version, these variants render the problem instances as images using simple geometrical shapes resulting in larger datasets and computational demand. Still we argue that the so generated tasks are not necessarily harder than the tasks of sraven, given that sraven models an additional difficulty of finding correspondences as detailed in Section 4.3 and its difficulty can be controlled parametrically; for instance by increasing the number of features.

### v4

@@ 3.1 Compositional generalization
We first investigate how well multi-head attention based transformer models are able to learn our fuzzy logic tasks in-context and compositionally generalize to held-out tasks. We vary the number of examples presented in context, as well as the fraction of combinations of fuzzy logic terms held-out from the training set. Figure 2C compares standard multi-head softmax attention and linear attention, as well as our HYLA model in terms of their coefficient of determination ($R^{2}$) on held-out tasks. We find that all models are able to solve the fuzzy logic task given sufficient examples are presented in context. As the fraction of held-out tasks increases, compositional generalization performance starts decreasing for all models. Interestingly, HYLA is less affected by this decline than softmax and linear attention. We hypothesize that the more expressive value network allows to more efficiently represent the nonlinear terms being composed and thereby strengthens the inductive bias towards compositionality. Consistent with this interpretation, in-distribution the models show comparably small differences in performance (see Figure A3B+D). In Appendix A.1 we conduct additional model ablations that separate the contributions of the RMSHead normalization and the nonlinearity added to the value network, showing that both components in combination are required to obtain the observed performance improvements.
Our compositional split only contains novel combinations of known terms. A priori, it might be possible that the system learns to compose the basic fuzzy logic operations in a way that allows it to generalize to any fuzzy logic function, including those that contain novel combinations of unknown terms. However, testing on functions obtained as novel combinations of $K$ unknown terms, we find that none of the models considered here is able to solve such a task (see Figure A3C).
@@ 3.2 Latent code structure
Next, we analyze the structure of the latent code for the different models solving the task. As suggested by Eq. 4, the attention scores for a given key-query pair across the head dimension configure the computation performed in the value network, and accordingly we would expect it to reflect the fuzzy logic function that needs to be implemented to correctly predict the output of the response token. To test this hypothesis, we collect the attention scores for the response token attending to itself, obtaining a vector of dimension $H$ for each layer and task. For this purpose, we use the tasks held-out from training, varying the response token while keeping inputs shown in-context fixed. In order to visualize the $H$-dimensional points we thereby obtain for each layer, we reduce the dimensionality using tSNE (Maaten & Hinton, 2008). The results of this analysis are shown in Figure 2D (for an extended figure showing all layers, see Figure A2). In line with our hypothesis, the latent codes for each model form clusters. The structure of these clusters is only partially explained by the specific output value of each function. Strikingly, coloring data points according to the fuzzy logic terms that make up each function reveals that clusters in the latent codes correspond to the terms. Indeed, it is possible to decode the terms underlying each function from the latent code using a logistic regression classifier which is trained to predict the function terms given the latents and is evaluated on held-out tasks as shown in Figure 2B.
@@
Figure 3: sraven. A Illustration of sraven task generation and the construction of the compositional generalization split. B Example problem instance illustrating a key challenge of the original Raven’s Progressive Matrices of finding correspondences (adapted from Carpenter et al. (1990)). When attempting to solve this instance, different hypotheses over which figural elements are governed by a consistent rule across rows are possible. This is akin to different orderings of the symbolic features.
Next, we introduce an abstract reasoning task based on Raven’s Progressive Matrices (Raven, 1962), we refer to as sraven. We develop this task to provide a challenging benchmark that probes symbolic reasoning capabilities while giving us fine-grained control over the constituent parts of each task in order to assess compositional abilities and analyze latent code structure.
4.1 Abstract reasoning based on symbolic Raven’s Progressive Matrices
Raven’s Progressive Matrices are a classic human intelligence test that probes the ability to induce abstract relations (Raven, 1962). It requires test subjects to generate and verify a potentially large number of hypotheses over rules that can parsimoniously explain a given problem instance (Carpenter et al., 1990). While there have been previous machine learning datasets inspired by Raven’s Progressive Matrices (Wang & Su, 2015; Zhang et al., 2019; Barrett et al., 2018; Hu et al., 2021), we seek to study a setting that is both symbolic, compositional and models a key difficulty of Raven tasks referred to as finding correspondences by Carpenter et al. (1990) which requires searching through a large number of possible hypotheses. We discuss existing Raven-based benchmarks in Section 6.
@@ 4.2 Task description
Figure 3A illustrates the generative process of sraven. For each task, a matrix of eight context panels is presented, and the goal is to predict the contents of the final response panel. In the original Raven’s Progressive Matrices Test, each panel typically displays a combination of geometrical shapes, as for instance shown in Figure 3B. For our symbolic version, each panel is defined as a tuple of $K$ integers that symbolically encode features (in our experiments we set $K=4$ unless stated otherwise). The $K$ features within each row evolve according to $K$ rules. Every rule is a function that takes two integers as input and outputs an integer. Figure 3A lists all rules, and Appendix B.3 describes each rule in detail. In order to maintain a finite set of possible symbols, we limit all tasks to contain only integers $\{0,1,2,\ldots,F-1\}$ (in our experiments, we set $F=8$). Arithmetic rules such as progression, addition or difference are therefore defined using modular arithmetic modulo $F$.
To create a task, we first sample $K$ out of the $R=8$ possible rules and then use each of the $K$ rules, to generate three sequences of length three given random inputs to each rule. As a result, we obtain nine panels, where each row can be explained by the same set of $K$ underlying rules. We use the first eight panels as context panels and present them sequentially to a sequence model, which needs to predict each feature of the response panel independently. A prediction is considered correct if and only if all subpredictions were correct. In the original Raven task, a number of possible answer panels are presented alongside the context panels and test subjects are asked to pick the one which contains the correct answer (Raven, 1962). Since in our case, the correct answer will consist of a combination of a finite set of known symbols, we can ask the agent solving the task to directly predict the missing symbols. Conveniently, this helps avoid potential shortcuts for solving a task that might arise from a biased generative process to create multiple-choice answers as a popular prior Raven-based benchmark has been found to be affected by (Zhang et al., 2019; Hu et al., 2021).
One of the key features of sraven is its compositional structure. Each task consists of a combination of a finite set of rule combinations. To systematically test whether a system trained on a subset of rule combinations can generalize to unseen combinations, we split all rule combinations into a train set and a test set. Unless noted otherwise, we hold out $25\%$ of all possible rule combinations for evaluation in our experiments. Note that permutations of rule combinations (e.g. AB and BA) are considered to be the same and will only appear in one of the two sets. This ensures that at test time, the agent is confronted with instances of tasks whose underlying combination of rules it has never encountered during training.
@@ 4.3 Finding correspondences
Making the task symbolic instead of presenting images of geometric shapes comes with the advantage of a smaller input space, allowing us to scale to larger model and data sizes in our experiments. However, it removes the visual representation learning step of discovering the right symbolic abstractions. We argue that while this is a difficult problem worth studying by itself, there is a major source of difficulty that arises even with access to symbolic abstractions. As Carpenter et al. (1990) argue in their theoretical account of the original Raven test, what makes many tasks difficult is the problem of finding correspondences of rules to feature dimensions.
This is best illustrated when considering the example in Figure 3B showing an adapted version of Figure 5 of Carpenter et al. (1990). When first attempting to solve this task, a test subject might hypothesize that there is a rule guiding the numerosity and orientations per shape. For example, one might try to find a rule that governs all curved shapes, one rule that governs all lines and one rule that governs all rectangles. Translated into a symbolic encoding, this is akin to ordering symbols by the shape and trying to find rules along each dimension in this ordering. However, this approach turns out to not provide a parsimonious answer. The correct answer can instead be obtained by grouping the objects by their orientations, i.e. by finding a rule that applies to all horizontal objects and a rule that applies to all vertical objects. In the symbolic encoding this new correspondence of features amounts to a new ordering of the same symbolic encoding where symbols are now ordered by orientation and one needs to find rules guiding the numerosity and shape per orientation.
To model the property of finding correspondences in sraven we similarly allow for permutations of the features. Specifically, we sample and apply a column-specific permutation of the features when generating each task, such that the entries of the tuples encoding each panel are permuted with a consistent permutation for a given column across all rows. This makes finding the correct solution more difficult, since a task can no longer be treated as $K$ independent tasks with a single feature and as a result the number of possible hypotheses grows significantly (see Appendix B.1 for more details). Note that in principle, a problem instance might have multiple valid but divergent answers, making the task ill-posed. In the settings considered here, this happens for less than half a percent of the problem instances. See Appendix B.2 for more details on handling such ambiguities.
@@ 4.4 Results
Figure 5: Latent code structure of sraven. A tSNE visualizations of the final layer latent codes for the final response token colored by the magnitude of the predicted target value. B Same as A but colored by the ground-truth rule the model needs to apply to generate the correct prediction. C OOD accuracy for varying numbers of attention heads. For a single head, the hypernetwork mechanism is absent, which hampers OOD generalization. D The difficulty of sraven can be parametrically controlled by varying, $K$, the number of features per panel . E Heatmap showing the pairwise cosine similarity between the average latent code of each rule for HYLA revealing how semantically related rules form clusters. For instance, rule F (addition) and G (difference) are implemented with a very similar latent code, indicating that the same code might be reused by flipping the sign of the operands. F Decoding performance of a logistic regression classifier trained to predict the ground-truth sraven rule based on the latent code at the final response token of training tasks and evaluated on unseen OOD tasks, revealing that the latent code is predictive of the implemented rule.
@@
We first evaluate to what extent transformers can compositionally generalize on sraven as we increase the number of problem instances trained on and scale up the models both in terms of width and depth. Figure 4 shows the results of these scaling experiments. Given sufficient data and model size, all model classes are able to solve 80% of problem instances created from tasks held-out during training. The inductive bias of making the value network more expressive by introducing a nonlinearity in HYLA seems to be beneficial for this task. Especially when trained on fewer problem instances and for small model sizes, it outperforms both linear and softmax attention.
Disrupting the hypernetwork mechanism hurts compositional generalization.
Varying the number of heads offers another way of manipulating the hypernetwork mechanism of multi-head attention. Specifically in the case of a single head the mechanism degenerates, only allowing the network to rescale the weights of the sole remaining value projection but no longer allowing it to compose multiple value projections, c.f. equation (4). In line with this argument, we observe a noticeable decrease in OOD performance in Figure 5C for the single head case.
@@
We next analyze the latent code for the different models solving the task. To this end, we collect the attention scores of the response token attending to itself. Following the perspective of attention as a hypernetwork, we expect latent codes corresponding to the same rule to form clusters. Indeed, Figure 5B reveals a strikingly structured code for the final layer across models, with clusters closely matching the rule to be inferred to correctly predict the output for each given task. In comparison, Figure 5A colors the points in latent space by the magnitude of the inputs the response token needs to attend to, to make the correct prediction. This tends to explain clusters more prominently in early layers, as shown in the appendix in Figure A6 and for a 16 layer softmax transformer in Figure A7.
To obtain a better understanding of the semantics of the latent code, we explore the pairwise cosine similarity between average latent codes for each rule across tasks for HYLA. Figure 5E shows that semantically related rules like sum/difference can form clusters, suggesting that latent codes might even be reused to solve related task operations. Finally, we train a logistic regression classifier to predict the ground-truth sraven rule based on the latent code at the final response token of in-distribution sraven tasks and evaluate it on unseen OOD tasks. This reveals that, especially for later layers, the latent code is strongly predictive of the subfunctions performed by the network.
Figure 6: Language modeling. Performance of autoregressively trained decoder-only transformer models with 50M parameters over the course of training on the C4 dataset with 130B tokens.
@@ 5 Language modeling
Multi-head attention has proven especially effective for language modeling. With language itself being highly compositional, it stands to question whether the implicit hypernetwork mechanism in multi-head attention might play a part in explaining its success. To shed light on this question, we conducted language modeling experiments and evaluate the performance of HYLA in comparison to linear and softmax attention. We train decoder-only transformer models with 50M parameters autoregressively for 130 Billion tokens on the C4 dataset (Raffel et al., 2020). We find that strengthening the hypernetwork mechanism via HYLA improves performance over linear attention and performs closely to softmax attention despite being a linear attention variant itself. This is noteworthy considering the importance of softmax in language modeling hypothesized to be due to its role in binding and associative recall problems on which linear attention methods typically struggle (Arora et al., 2023; Olsson et al., 2022; Schlag et al., 2021). The fact that reinforcing the hypernetwork mechanism implicit in multi-head attention as done by HYLA helps close the gap of linear attention to softmax attention suggests that the hypernetwork mechanism might be of practical relevance for understanding large-scale transformer models.
@@ 6 Related work
Prior work that has studied the question of why the attention mechanism benefits from multiple heads has counterintuitively found that it is possible to prune almost all but one head in some layers after training, sacrificing only relatively small amounts of performance (Voita et al., 2019; Michel et al., 2019). It has therefore been speculated that equipping the attention mechanism with multiple heads primarily aids stability during training (Liu et al., 2021). The hypernetwork perspective of attention offers an alternative account, suggesting that multiple heads create a way of configuring compositional computations in the value network. Seen from this perspective, the observation of singular heads dominating could point to a connection to the phenomenon of module collapse often observed when training modular systems in practice (Shazeer et al., 2017; Kirsch et al., 2018; Rosenbaum et al., 2018; Mittal et al., 2022).
@@
Consistent with our observation on scaling model size on the sraven task, Hosseini et al. (2022) find that as pretrained large language models are scaled, their ability to compositionally generalize on in-context semantic parsing tasks improves. Similarly, outside the in-context learning regime, Furrer et al. (2021) have demonstrated that pretrained large language models outperform many architectures specialized towards compositional generalization. Still, to what extent the ability for compositional generalization extends beyond the training distribution remains openly debated (Srivastava et al., 2023; Press et al., 2023; Dziri et al., 2023).
@@
A number of Raven-inspired tasks have been introduced to assess abstract reasoning capabilities of neural networks (Wang & Su, 2015; Zhang et al., 2019; Barrett et al., 2018; Hu et al., 2021). Common to all these variants, including our version is an underlying generative task model based on a finite number of possible rules that are applied on one or more of the feature dimensions for each panel. Different from our version, these variants render the problem instances as images using simple geometrical shapes resulting in larger datasets and computational demand. Still, we argue that the so generated tasks are not necessarily harder than the tasks of sraven, given that sraven models an additional difficulty of finding correspondences as detailed in Section 4.3 and its difficulty can be controlled parametrically; for instance by increasing the number of features.